OpenStudy (timaashorty):

For questions 4 and 5, what values for θ (0 ≤ θ ≤ 2 pi ) satisfy the equation? 4. 4cos θ + 1 = 2cos θ a. 2pi/3, 3pi/3 b. pi/6, 5pi/6 c. 7pi/6,11pi/6 d. pi/3 6pi/3

OpenStudy (timaashorty):

5. cos θ - tan θ cos θ = 0 a. 0, pi/4, pi, 5pi/4 b. pi/4, 5pi/4 c. pi/2, 3pi/4, 3pi/2, 7pi/4 d. pi/2, 7pi/6, 3pi/2, 11pi/6

hero (hero):

Solve for theta for both equations

OpenStudy (timaashorty):

1. theta = 2pi/3 +p 2pin,4pi/3 += 2pin

OpenStudy (timaashorty):

So for the first one is A ?

OpenStudy (timaashorty):

\[\frac{ 2 \pi }{ 3 }, \frac{ 3\pi }{ 3 }\]

OpenStudy (timaashorty):

2. \[\theta = \frac{ \pi }{ 2 }\pm 2\pi n, \frac{ 3\pi }{ n }\pm 2\pi n, \frac{ \pi }{ 4 }\pm \pi n,\frac{ 5\pi }{ 4}\pm \pi n\]

OpenStudy (timaashorty):

@Hero ?

OpenStudy (timaashorty):

---- 3pi/2 not n ----

OpenStudy (timaashorty):

I'm confused on the second one.. I think it's either B or C.

OpenStudy (timaashorty):

...?

hero (hero):

There seems to be a typo for one of your answer choices.

OpenStudy (timaashorty):

I got my answers already [: thanks for giving me a start!